OpenStudy (omgthatgirl):

Find the first three output values of the fractal generating function f(z)=z^2+2+i a. 2+i, 2+5i, -2595+205i b. 2+i, 5+5i, -2595+205i c. 2+i, 5+5i, 2+51i d. 5+5i, 2+51i, -2595+205i

OpenStudy (perl):

We can use the multiple choices to tell us the initial value. Let z_0 = 2 + i . now find f(2 + i )

OpenStudy (omgthatgirl):

how do i do that? sorry i dnt know im prettty slow

OpenStudy (omgthatgirl):

oh wow

OpenStudy (omgthatgirl):

ok

OpenStudy (omgthatgirl):

whats next

OpenStudy (perl):

Suppose the first term is 2 + i, then the second term is f(2 + i ) $$ \Large { f(2+i) = (2+i)^2 + 2 + i \\= (2+i)(2+i) + 2 + i \\=2^2 + 2i + 2i + i^2 + 2 + i \\ = 4 + 4i + (-1) + 2 + i \\ = 5 + 5i } $$

OpenStudy (perl):

now take this value and plug into f(z)